OpenStudy (anonymous):

Which expression is equivalent to (7^3)^-2?

Parth (parthkohli):

You multiply the powers.

Parth (parthkohli):

\( \color{Black}{\Rightarrow (7^3)^{-2} = 7^{-6}}\)

OpenStudy (anonymous):

ok the answers we can choose from are 7 1/7 1 over six 7's or 1 over seven 6's

OpenStudy (apoorvk):

\[\large (a^m)^n = a^{mn}\]

OpenStudy (anonymous):

ok i get 7^6 but how do i get it to the correct answer 7 1/7 1 over six 7's or 1 over seven 6's

OpenStudy (apoorvk):

okay, one more thing, \[\large a^{-b}=\frac 1 {a^b}\] so, what's \(7^{-6}\) then?

OpenStudy (apoorvk):

@brett183 ??

OpenStudy (anonymous):

i am thinking it is the 1 over the six 7's but not sure

OpenStudy (theviper):

\[\frac{1}{7}^6\]

OpenStudy (apoorvk):

yeah, you're thinking the right way, correct-o!! @brett183

OpenStudy (apoorvk):

@TheViper Uh-oh slight mistake in your post

OpenStudy (theviper):

yeah

OpenStudy (anonymous):

what

OpenStudy (anonymous):

oh never mind

OpenStudy (anonymous):

so i multiply the powers then put a one over the 7 with the power of 6 and that tells me how many 7's i need right

OpenStudy (apoorvk):

exactly!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i have another question do you have time

OpenStudy (anonymous):

Explain how you would find the product of 20^12 and 20^8. Include either the exponent rule or the expanded method in your explanation. Give your answer in exponential form.

OpenStudy (apoorvk):

Hmm, use this funda: \[\large a^m \times a^n = a^{m+n}\] Can you use this?

OpenStudy (anonymous):

i dont know it is a esay question. and i am taking a online class.

OpenStudy (apoorvk):

yeah so use this formula, the 'a' in your case is 20, the 'm' is 12, and the 'n' is '8', so, \(20^{12} \times 20^8 = ??\) just use the formula i posted above

OpenStudy (anonymous):

ok so would i do the same for this one. A kilobyte is 210. A megabyte is 220 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form.

OpenStudy (anonymous):

but subtract

OpenStudy (apoorvk):

you mean \(2^{10}\) and \(2^{20}\) right? yeah, so you would divide, and this time, we subtract the exponents! \[\large \frac{a^m}{a^n}= a^{m-n}\]

OpenStudy (anonymous):

yeap thats what i mean. thank you

OpenStudy (apoorvk):

Hey please post new questions in fresh threads after 'closing' this one. Thanks :)

OpenStudy (anonymous):

hey i have one that i got wrong but i didnt ask you this one. simplify 15^-4 over 15^5. i answered 15^1

OpenStudy (anonymous):

sorry didnt see your post. before i sent the question

OpenStudy (apoorvk):

umm, when it's "over", means you're dividing, so you subtract the exponents. so, \[\large \frac{15^{-4}}{15^5} = 15^{-5-4} = 15^{-9}\]

OpenStudy (apoorvk):

lol no worries ;)

OpenStudy (anonymous):

ok thank u

OpenStudy (apoorvk):

anytime!